the sum of three times a number and 7 more than the number is the same as the difference between -11 and twice a number. what is the number?
Answer by edjones(8007)

Let x be the number.
3x+x+7=-11-2x
4x+7=-11-2x
6x=-18
x=-3
.
Ed